OpenStudy (anonymous):

Write the first five terms of the sequence defined recursively. Use the pattern to write the nth term of the sequence as a function of n. (Assume that n begins with 1.) a[1]=5,a[k+1]=-a[k]

OpenStudy (anonymous):

a[1] = 5, a[2] = -5, a[3] = 5, a[4] = -5, a[5] = 5

OpenStudy (anonymous):

ok but my answer choices are: 1.a[n]=5n 2.a[n]=5(-1)^n 3.a[n]=(-5)^n 4.a[n]=(-5)^(n-1) 5.a[n]=5(-1)^(n-1)

OpenStudy (anonymous):

SO where do I go from there?